# Enabling Fast and Accurate Neutral Atom Readout through Image Denoising

### v1

@@ 5.3. Pipelining classification across cycles
In current approaches, classification is performed immediately after image acquisition, with inference latency accumulating at every QEC cycle as in Figure.10(a). While negligible for small systems, this overhead becomes significant as arrays scale to hundreds of qubits. Baseline methods typically employ both single-site and multi-site CNN classifiers with classification taking upto s per site, and s for a array, respectively, almost an order of magnitude slower for multi-site. Similar scaling to larger arrays quickly drives classification into the multi-millisecond regime, creating a major throughput bottleneck.
Our approach mitigates this by relying on lightweight single-site classifiers such as FNNs and MF+NNs runs parallel across sites and achieve up to faster inference than CNN(Array). Multi-site correlations are instead captured in the denoising stage, allowing classification to remain compact and scalable. Moreover, we propose pipelining classification with image acquisition to effectively hide its latency, eliminating it as a limiting factor for large-scale QEC.
In existing small-scale NA readout systems, classification is performed immediately after image acquisition to trigger reset and atom reloading, making it a per-cycle bottleneck. Recent works suggest the reset(Gehér et al., 2025) and atom loss(Baranes et al., 2025) can be tolerated. As chances of atom loss is rare and does not introduce spurious correlations, reloading can instead be deferred by one round, allowing classification to overlap with the next measurement as shown in Fig. 11(b). This reduces idle time and mitigates idling-induced errors.
With unpipelined design, time to execute QEC rounds:
@@ 6. Methodology
Quantum hardware and dataset. We obtained datasets with the readout images from the 9-qubit neutral atom system mentioned in (Phuttitarn et al., 2024) with two different atom spacings as and resulting in and readout images respectively. The qubit atom array is illuminated for a variable readout duration and the scatter photons are collected using EMCCD camera. A total of 14 different readout durations ranging from 15ms to 100ms are used to obtain long-exposure images. Additional path is setup to attenuate the signal by resulting in short-exposure images with effective readout durations ranging from 1.5ms to 10ms.
@@
Simulation framework. To evaluate the effect of readout errors on logical error rates of surface codes and BB codes using Stim(Gidney, 2021) stabilizer simulator with PyMatching and BP-OSD-CS10 as the respective decoders.
Hardware. We evaluate the inference latency on NVIDIA GeForce RTX 2080 Ti and Intel(R) Xeon(R) Silver 4216 CPU.
@@ 7. Evaluations
Evaluation metrics. We evaluate GANDALF at three levels: (1) State Classification: averaged-site inaccuracy and relative infidelity reduction () over the baseline, including post-selection filtering to capture accuracy–latency tradeoffs. (2) System level: logical error rate (LER) for surface and bicycle bivariate codes, accounting for idling errors.(3) QEC cycle latency: execution-time reduction for QEC codes.
@@ 7.1. Impact on readout accuracy and readout duration
Figure 12 shows the improvement in readout inaccuracy as a function of readout duration for both m and m atom spacings. GANDALF delivers substantial gains at short exposure times, where SNR is lowest. Our method achieves reduction in readout error compared to the baseline CNN(Site) at 1.5ms. Furthermore, our method sustains below 1% inaccuracy up to 2.9ms for m and 3.2ms for m, compared to 3.6ms and 4ms for the baseline, respectively. On average, relative infidelity factor, i.e, , where denotes the classification inaccuracy, improves by 64% for atom spacing m and 53% for m lattices.
The dataset used in this work is imperfect, but higher-quality data is expected in the near future. To estimate the best-case improvements achievable with denoising, we construct a confidence-filtered dataset using Gaussian Mixture Model (GMM)–based confidence filtering. Instead of relying solely on raw classifier outputs, we fit a two-component GMM to the distribution of decision scores, separating high-confidence bright and dark populations from ambiguous cases as shown in Figure 13(b). By post-selecting only the confident subsets, we maintain sub-1% readout inaccuracy even at the shortest duration of 1.5ms as shown in Figure 13(a). This demonstrates the potential to reduce readout times close to sub-millisecond regime for Cs arrays aiding GANDALF with confidence-aware post-selection.
@@ 7.2. Logical error rate
To quantify the end-to-end impact of readout fidelity and latency on fault-tolerant performance, we evaluate logical error rates (LER) using both bivariate bicycle (BB) codes and surface codes (SC). Idling errors during measurement were modeled through the Pauli twirling approximation (Tomita and Svore, 2014) characterized by and times, allowing us to capture the tradeoff between longer readout times with improved accuracy but increased idling error against shorter readout times with reduced idling error but lower accuracy.
Figure 14 indicates an intermediate readout duration of approximately 4ms achieves the best overall LER among the readout durations for 9um, similar trends are observed for 5um. GANDALF-enhanced FNN achieves reduction in LER compared to CNN(Site) baseline at 4ms and upto reduction at 1.5ms for BB codes, highlighting the benefits of better readout accuracy at shorter readout duration.
@@ 7.3. Reducing QEC cycle latency
Pipelining overlaps denoising and classification with subsequent acquisitions, so their cost is paid only once at the final round. As shown in Fig. 15, this reduces QEC execution time by over 10% () across 100 rounds, with the largest gains at short readout durations where classification overhead dominates. Incorporating GANDALF further improves performance, achieving up to speedup over baseline.
At shortest readout time of , the execution time for 100 rounds decreases from with the baseline to with GANDALF and pipelining. The advantage grows with number of QEC rounds, as in conventional pipelines, classification adds constant latency with rounds, while the pipelined scheme hids it. As a result, classification never re-emerges as a bottleneck, when less than readout time, ensuring scalability of real-time QEC to large neutral atom arrays.
@@ 8. Conclusions
We present GANDALF, a system architecture that optimizes neutral atom readout by addressing fidelity and latency, the main bottlenecks in scalable quantum error correction. Our framework uses explicit denoising with image translation to reconstruct clear signals from short, low-photon measurements, enabling reliable classification at up to shorter readout times. Combined with lightweight classifiers and a pipelined readout design, GANDALF reduces logical error rate by for BB code and by for Surface code and enables lower QEC cycle time compared to state-of-the-art CNN-based readout for neutral atom arrays. Fully convolutional networks used in our design enable scaling to very large arrays, with batching and parallelism achieving millisecond-level throughput, together establishing a practical path to high-fidelity, low-latency, and scalable readout for large-scale neutral atom processors.

### v2

@@ 5.3. Pipelining classification across cycles
In current approaches, classification is performed immediately after image acquisition, with inference latency accumulating at every QEC cycle as in Figure 11(a) and without GANDALF, the acquisition time is significantly larger compared to when denoising is enabled as shown in Figure 11(b). While the sequential execution overhead is negligible for small systems, it becomes significant as arrays scale to hundreds of qubits. Baseline methods typically employ both single-site and multi-site CNN classifiers with classification taking up to s per site, and s for a array, respectively, almost an order of magnitude slower for multi-site. Similar scaling to larger arrays quickly drives classification into the multi-millisecond regime, creating a major throughput bottleneck.
Our approach mitigates this by relying on lightweight single-site classifiers such as eedforward neural networks and matched filter with neural networks (MF+NN) that run in parallel across sites and achieve up to faster inference than CNN (Array). Multi-site correlations are instead captured in the denoising stage, allowing classification to remain compact and scalable. Moreover, we propose pipelining classification with image acquisition to effectively hide its latency, eliminating it as a limiting factor for large-scale QEC.
In existing small-scale neutral atom readout systems, classification is performed immediately after image acquisition to trigger reset and atom reloading, making it a per-cycle bottleneck. Recent works suggest that reset(Geher2025) and atom loss(baranes2025leveragingatomlosserrors) can be tolerated. As the chance of atom loss is rare and does not introduce spurious correlations, reloading can instead be deferred by one round, allowing classification to overlap with the next measurement as shown in Fig. 11(b). This reduces idle time and mitigates idling-induced errors.
With an unpipelined design, the time to execute rounds of syndrome extraction for QEC is,
@@ 6. Methodology
Quantum hardware and dataset. We obtained datasets with readout images from the atom array of with 9-qubit neutral atom system as mentioned in (phuttitarn2023enhanced) with two different spacings between atoms of and resulting in and pixel images respectively. The qubit atom array is illuminated for a variable readout duration and the scattered photons are collected using an EMCCD camera. A total of 14 different readout durations ranging from to are used to obtain single-site long-exposure images for each duration. For each long exposure image, the dataset has corresponding short-exposure images obtained by attenuating the signal by , resulting in effective readout durations ranging from to .
@@
Simulation framework. The effect of readout errors on logical error rates of surface codes and BB codes is evaluated using the Stim(gidney2021stim) stabilizer simulator with PyMatching and BP-OSD-CS10 as the respective decoders.
Hardware. We evaluate our models on NVIDIA® GeForce RTX™ 2080 Ti GPU and Intel® Xeon™ Silver 4216 CPU.
@@ 7. Evaluations
Evaluation metrics. We evaluate our method, GANDALF, at three levels: (1) State Classification: averaged-site inaccuracy and relative infidelity reduction () over the baseline, including post-selection filtering to capture accuracy–latency tradeoffs. (2) System level: Logical Error Rate (LER) for surface code and Bivariate Bicycle (BB) code, accounting for idling errors. (3) QEC cycle latency: execution-time reduction of repeated syndrome extraction for QEC codes.
@@ 7.1. Impact on readout accuracy and readout duration
Figure 12 shows the improvement in readout inaccuracy as a function of readout duration for both m and m atom spacings. GANDALF delivers substantial gains at short exposure times, where SNR is lowest. Our method achieves reduction in readout error compared to the baseline CNN(Site) at . Furthermore, our method sustains below 1% inaccuracy up to for m and for m, compared to and for the baseline, respectively. On average, relative infidelity factor, i.e, , where denotes the classification inaccuracy, improves by 64% for atom spacing m and 53% for the m lattices.
The dataset used in this work is imperfect, but higher-quality data is expected in the near future. To estimate the best-case improvements achievable with denoising, we construct a confidence-filtered dataset using Gaussian Mixture Model (GMM)–based confidence filtering. Instead of relying solely on raw classifier outputs, we fit a two-component GMM to the distribution of decision scores, separating high-confidence bright and dark populations from ambiguous cases as shown in Figure 13(b). By post-selecting only the confident subsets, we maintain sub-1% readout inaccuracy even at the shortest duration of as shown in Figure 13(a). This demonstrates the potential to reduce readout times close to sub-millisecond regime for Cs arrays aiding GANDALF with confidence-aware post-selection.
@@ 7.2. Logical error rate
To quantify the end-to-end impact of readout fidelity and latency on fault-tolerant performance, we evaluate logical error rates (LER) using both bivariate bicycle (BB) codes and surface codes (SC). Idling errors during measurement were modeled through the Pauli twirling approximation (Tomita2014) characterized by and times, allowing us to capture the tradeoff between longer readout times with improved accuracy but increased idling error against shorter readout times with reduced idling error but lower accuracy.
Figure 14 indicates an intermediate readout duration of approximately 4ms achieves the best overall LER among the readout durations for spacing, similar trends are observed for spacing. GANDALF-enhanced FNN achieves reduction in LER compared to the CNN(Site) baseline at and upto reduction at for BB codes, highlighting the benefits of better readout accuracy at shorter readout duration.
@@ 7.3. Reducing QEC cycle latency
Pipelining overlaps denoising and classification with subsequent acquisitions, so their cost is paid only once at the final round. As shown in Fig. 15, this reduces QEC execution time by over 10% () across 100 rounds, with the largest gains at short readout times, where classification dominates the cost. Incorporating GANDALF further reduces QEC cycle time, achieving up to reduction in the pipelined baseline.
At the shortest readout time of , the execution time for 100 rounds decreases from with the baseline to with GANDALF and pipelining. The advantage grows with number of QEC rounds as conventional pipelines adds constant classification latency per round. While the overlapping pipelined scheme removes this latency from critical path.
@@ 8. Conclusions
We present GANDALF, a system architecture that optimizes neutral atom readout by addressing fidelity and latency, the main bottlenecks in scalable quantum error correction. Our framework uses explicit denoising with image translation to reconstruct clear signals from short, low-photon measurements, enabling reliable classification at up to shorter readout times. Combined with lightweight classifiers and a pipelined readout design, GANDALF reduces the logical error rate up to for BB code and up to for Surface code and enables lower QEC cycle time compared to state-of-the-art CNN-based readout for neutral atom arrays. Fully convolutional networks used in our design enable scaling to very large arrays, with batching and parallelism achieving millisecond-level throughput, together establishing a practical path to high-fidelity, low-latency, and scalable readout for large-scale neutral atom processors.